OpenStudy (anonymous):

If f(x) = \frac { 4 } {x^2}, find f'( 3 ).

OpenStudy (anonymous):

\[\large{ f(x) = \frac { 4 } {x^2}=4x^{-2}\\ f'(x)=4(-2)x^{-3}=-\frac{8}{x^3},f'(3)=?}\]

OpenStudy (cggurumanjunath):

put x=3 in (-8/(x^3)) as shown above by @Jonask ;@kquach1.